please help answer:)a reactangles area is three -fourths,if the width is one half,what is the length? i tried three-fourths divided by one half but i dont thing that is right
Answer by rothauserc(4718)

we are given
a reactangle's area is three -fourths,if the width is one half,what is the length?
area of rectangle = length * width
so we have
3/4 = 1/2 x
multiply both sides of = by 2
6/4 = x
and x = 3/2